[figure: 9.7 cm]

ON THE MAGNETIC BEHAVIOUR OF GdCo_{2−x}Cu_{x} COMPOUNDS

R. Tetean^{a}, E. Burzo^{a}, L. Chioncel^{b}, T. Crainic^{a} a Faculty of Physics, Babes-Bolyai University, RO-3400, Cluj-Napoca, Romania b University of Nijmegen, 6500-GL Nijmegen, The Netherlands

The GdCo_{2−x}Cu_{x} compounds crystallize in a cubic C15 type structure for x≤0.4. Magnetic measurements were performed in the temperature range 4.2-900 K and external fields up to 9 T. Band structure calculations were carried out by using the ab initio tight binding linear muffin tin orbitals method in the atomic sphere approximation.The compounds are ferrimagnetically ordered. The mean cobalt magnetic moments determined at 4.2 K are little dependent on composition. The saturation moments per formula unit, at 4.2 K, are in rather good agreement with those obtained from band structure calculations. The Gd 5d band polarizations were analyzed. The contribution from 4f-5d, M_{5d}(f), local exchange and 5d-3d short range exchange interactions, M_{5d}(0), were determined. The M_{5d}(0) values are proportional to cobalt magnetizations. The paramagnetic behaviour was also analyzed. Finally, the exchange interactions were discussed in correlation with band structure calculations.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : R. Tetean

Address for correspondence : Faculty of Physics Babes-Bolyai University RO-3400 Cluj-Napoca ROMANIA

Email address : rote@phys.ubbcluj.ro